OpenStudy (anonymous):

Solve for x using mental math 7(x - 3) + 3(4 - x) = -8

OpenStudy (anonymous):

Just add up the terms dear 7x - 21 + 12 - 3x = -8 Simplify this and ur answer is on the way

OpenStudy (anonymous):

then what

OpenStudy (anonymous):

U will get x value it is like 4x - 9 = -8 => 4x = -8 + 9 =1 =>x = 1/4

OpenStudy (the_fizicx99):

@ganibl is right. You get 1/4 for x

OpenStudy (raffle_snaffle):

distribute the #'s through then solve for x

OpenStudy (anonymous):

Ok. thanks guys. I suck at mental math and memorising numbers

OpenStudy (the_fizicx99):

xD I'm sure after a while you'll get better. I did 6 and 7th grade math problems in my head.... bad idea, barely remember anything but it was right at that moment lol :/